OpenStudy (cici_lyn4):

Write an equation of a line that passes through the point (3, 2) and is parallel to the line y = 3x − 4. y = 3x + 7 y = 3x − 7 y = 1 over 3x + 2 y = 1 over 3x − 2

OpenStudy (sshayer):

any equation parallel to ax+by=c is ax+by=k substitute the point and find k finally substitute the value of k

OpenStudy (queenie):

y = 3x - 4...the slope on this equation is 3....a slope on a parallel equation will be the same. Now, we use y = mx + b slope(m) = 3 (3,2)....x = 3 and y = 2 now we sub...we already have out points x and y....and we have our slope (m)...all we need is b, the y intercept...so we sub and solve for b. y = mx + b 2 = 3(3) + b 2 = 9 + b 2 - 9 = b -7 = b so ur parallel equation is : y = 3x - 7